The lucid and lucmath packages^{∗}

Sebastian Rahtz

1997/01/31

1 Packages

1.1 Lucid Package

Set up text fonts for Adobe Lucida.

1 ⟨∗lucid⟩

 2 \renewcommand{\sfdefault}{plcs}
 3 \renewcommand{\rmdefault}{plc}
 4 \renewcommand{\ttdefault}{pcr}

5 ⟨/lucid⟩

1.2 Lucmath Package

Set up math fonts for Adobe Lucida.

6 ⟨∗lucmath⟩

 7 \DeclareSymbolFont{operators}{\encodingdefault}{\rmdefault}{m}{n}
 8 \DeclareSymbolFont{letters}{OML}{plcm}{m}{it}
 9 \DeclareSymbolFont{symbols}{OMS}{plcy}{m}{n}
10 \DeclareSymbolFont{largesymbols}{OMX}{plcv}{m}{n}
11 \SetSymbolFont{letters}{bold}{OML}{plcm}{m}{it}
12 \SetSymbolFont{operators}{bold}{\encodingdefault}{\rmdefault}{b}{n}
13 \SetSymbolFont{operators}{normal}{\encodingdefault}{\rmdefault}{m}{n}

reset the text in math macros

14 \DeclareMathAlphabet      {\mathbf}{\encodingdefault}{\rmdefault}{b}{n}
15 \DeclareMathAlphabet      {\mathrm}{\encodingdefault}{\rmdefault}{m}{n}
16 \DeclareMathAlphabet      {\mathsf}{\encodingdefault}{\sfdefault}{m}{n}
17 \DeclareMathAlphabet      {\mathit}{\encodingdefault}{\rmdefault}{m}{it}
18 \DeclareMathAlphabet      {\mathtt}{\encodingdefault}{\ttdefault}{m}{n}
19 \SetMathAlphabet{\mathbf}{bold}{\encodingdefault}{\rmdefault}{b}{n}
20 \SetMathAlphabet{\mathsf}{bold}{\encodingdefault}{\sfdefault}{b}{n}
21 \SetMathAlphabet{\mathrm}{bold}{\encodingdefault}{\rmdefault}{b}{n}
22 \SetMathAlphabet{\mathit}{bold}{\encodingdefault}{\rmdefault}{b}{it}
23 \SetMathAlphabet{\mathtt}{bold}{\encodingdefault}{\ttdefault}{b}{n}

24 ⟨/lucmath⟩

∗This file has version number v4.0, last revised 1997/01/31.

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones underlined to the code line of the definition, the rest to the code lines where the entry is used.

D

\DeclareMathAlphabet

. . . . . . . . . 14–18

\DeclareSymbolFont .

. . . . . . . . . . 7–10

E

\encodingdefault . .

. . . . . . . . 7, 12–23

M

\mathbf . . . . . . . . 14, 19 \mathit . . . . . . . . 17, 22 \mathrm . . . . . . . . 15, 21 \mathsf . . . . . . . . 16, 20 \mathtt . . . . . . . . 18, 23

R

\renewcommand . . . . . 2–4

                          \rmdefault 3, 7, 12–

15, 17, 19, 21, 22

S

                          \SetMathAlphabet 19–23
                          \SetSymbolFont . . 11–13
                          \sfdefault . . . 2, 16, 20

T

                          \ttdefault . . . 4, 18, 23